Sketch the function f(x)=2/x on the interval 1/2<= x <= 4
Answer by Edwin McCravy(20060)

Sketch the function f(x)=2/x on the interval 1/2<= x <= 4
f(0.5) = 2/(0.5) = 4
f(1.0) = 2/(1.0) = 2
f(1.5) = 2/(1.5) = 1.333
f(2.0) = 2/(2.0) = 1
f(2.5) = 2/(2.5) = 0.8
f(3.0) = 2/3 =     0.667 
f(3.5) = 2/(3.5) = 0.541
f(4.0) = 2/4 =     0.5

plot those points and sketch a smooth curve through them:




Edwin